OpenStudy (safa102):

What is the slope-intercept form of the function that contains the points (3,4) and has a slope of 2?

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (skullpatrol):

Recall the slope-intercept form: y=mx+b.

OpenStudy (skullpatrol):

y=mx+b m=2 x=3 y=4 then solve for b

OpenStudy (skullpatrol):

$$\Huge y=mx+b$$ $$\Huge4=2(3)+b$$

OpenStudy (safa102):

I STILL DON'T GET IT?

OpenStudy (triciaal):

the slope is the same for any 2 points on that line. you have 2 points (3,4) and (x, y) and you know the slope is 2 the slope, m is the change in y divided by the corresponding change in x rearrange the equation to be in the slope-intercept format y = m x + b

OpenStudy (safa102):

ohhh thanks so much

OpenStudy (safa102):

i get it now

OpenStudy (triciaal):